OpenStudy (anonymous):

Write this formula in center form: x^2 - 14x + 36y^2 - 216g +337 = 0

OpenStudy (anonymous):

Can you reprase the question? I am a bit unclear as to what you mean, because I have either never heard the term or have forgotten due to a long period without physical school

OpenStudy (anonymous):

It's a question that my pre calc teacher gave me for homework. We've been learning on how to write questions in center form for an ellipse. She asked up for this question to take this standard form and turn it into center form for an ellipse which is \[((x-h)^{2}/ a ^{2}) + ((y-k)^{2}/ b ^{2})\]

OpenStudy (anonymous):

Ahh ok, precalc is not my forte, youll have to ask some one else this time :) sorry

OpenStudy (anonymous):

thanks anyway! =)

OpenStudy (campbell_st):

complete the square for terms in x and then again for terms in y so x^2 -14x +49 36y^2 -216y + 324 = -337 + 49 + 324 (x - 7)^2 +(6y - 18)^2= 6^2 the centre is (7, 3) and radius = 6 units

OpenStudy (anonymous):

how did you get 36y^2 - 216y + 324? where did the 324 come from? dont you have to half 216 and square it?

OpenStudy (campbell_st):

given (a + b)^2 the middle term is double the product of a and b so start my halving -216y. This gives -108y divide by 6y (since(6y)^2 = 36y^2) leaves -18... so the b value is -18...

OpenStudy (anonymous):

oooo okay! thank you!